A grandfather is ten times older than his granddaughter. He is also 54 years older than her. Find their present ages. 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let the age of granddaughter be x and grandfather be 10x.
Also, he is 54 years older than her.
According to the question,
10x = x + 54
⇒10x-x=54
⇒9x=54
⇒x=6
Age of grandfather = 10x = 10×6 = 60 years.
Age of granddaughter = x = 6 years.

Thanks